# Diffusion MRI with Machine Learning

### v1

@@ II-D Delineation of white matter tracts
  • In addition to the streamline similarity measures, information such as homology between the brain hemispheres, spatial and shape priors, and population averages (atlases) can be used to improve the accuracy of tractography-based methods. However, all tractography-based methods are inherently limited by the errors in streamline tracing [19, 191]. Moreover, they typically involve several processing steps such as streamline tracing, white matter segmentation, gray matter parcellation, and clustering or similar computations to extract individual tracts. Each of these computations may introduce its own errors that depend on method settings, and it is difficult to jointly optimize all these processing steps. Moreover, some of these computations, e.g., tractography and streamline clustering, can require long computation times.
  • To avoid the errors and limitations mentioned above, several studies have proposed to segment the tracts on diffusion tensor or fiber orientation images, thereby avoiding the tractography. Some of the segmentation methods that have been explored in the past include Markov Random Fields [218], k-nearest neighbors technique [219], level-set methods [220, 221, 222], template/atlas-based techniques [223, 191], and more recently deep learning [27, 224].
  • Several works have successfully used deep learning methods to segment white matter tracts based on local fiber orientations. Wasserthal et al. [37] proposed to represent overlapping/crossing tracts using a simple but powerful concept that they named tract orientation maps (TOMs). Each tract is represented by a separate TOM, where each voxel contains a vector representing the local orientation of that tract. If a tract does not cross a certain voxel, the value of that voxel in the TOM for that tract will be a vector of zeros. In their framework, a standard method such as CSD is applied to compute the FOD and the local peaks of the FOD are extracted to build the TOMs. These peak orientations are fed into an FCN, which is trained to compute the TOMs. These TOMs can be used as a prior or as the direct input for bundle-specific tractography. The authors claim that their method is a viable solution to the problem of excessive false positive rate of standard tractography techniques. They support this claim by demonstrating that their method leads to superior tract bundle reconstruction results compared with several state of the art methods such as TractQuerier [209], RecoBundles [199], and WhiteMatterAnalysis (WMA) [182]. In subsequent works, they build upon this method and extend it in a few important directions. In TractSeg [27], they develop an FCN-based method to map the TOMs to the probability maps depicting the extent of each white matter tract. To enable the method to run on limited GPU memory while working on full-resolution images and predicting the probability map for the full set of 72 tracts, their model is a 2D FCN that works on axial, coronal, and sagittal slices separately. TractSeg achieves a mean DSC of 0.84. On all tracts except for commissure anterior and fornix it achieves a mean DSC of higher than 0.75. While the tract segmentation masks generated by TractSeg can be used for accurate bundle-specific tractography [27], a more complete method is presented in a subsequent work by the same authors [225], where a separate deep learning model segments the start and end regions of each tract.
  • A number of studies have built upon the methods proposed by Wasserthal et al. [37, 225]. A representation similar to TOM was proposed in HAMLET [163], where the model output is a tensor field that indicates the presence and local orientation of a tract. Li et al. [226, 197] trained a CNN to segment 25 white matter tracts directly from diffusion tensor images. Their extensive experiments showed high segmentation accuracy, low scan-rescan variability in tract-specific FA analysis, good generalizability to cross-center data, and successful application in a tract-specific population study on dementia patients. They showed that their method was more accurate than tractography-based techniques, while being orders of magnitude faster. They also found that adding anatomical (T1) images to the diffusion tensor image as input did not improve the segmentation accuracy. Moreover, using the directions of the three major peaks led to lower accuracy than using the diffusion tensor [197]. Liu et al. follow an approach similar to TractSeg but propose to model the correlation between tract labels [186]. Specifically, they introduce an auxiliary neural network that maps the native tract labels into a lower-dimensional label space. The main segmentation network predicts this lower-dimensional label, which can be used to recover the tract presence in each voxel. The authors argue that this makes the task of tract segmentation simpler to learn and show that it can improve the segmentation of certain tracts such as the fornix that are especially difficult to segment by TractSeg. Similarly, Lu et al. [193] follow the general approach originally proposed by TractSeg, but propose to perform self-supervised pre-training to reduce the required amount of labeled data. They propose two pretext tasks: estimating the tractography-generated streamline density and estimating tract segmentations computed with a multi-atlas method. An alternative method based on transfer learning was proposed by the same authors in [227]. These works have shown that as few as five labeled training images may be sufficient to train a CNN for segmenting certain white matter tracts. To reduce the required number of labeled training images, one study proposed a deep learning registration method to align the labeled data to the space of unlabeled images, thereby synthesizing a much larger labeled dataset [198]. Experiments showed that even a single labeled image was sufficient to train an accurate deep learning model.
  • In order to determine the orientation of major fascicles, previous works have relied on computation of the diffusion tensor or FOD. However, none of these intermediate computations have an unambiguous biophysical meaning and they entail unavoidable estimation errors. Moreover, the intermediate computations for most existing methods assume availability of dense multi-shell measurements, which are not acquired in many clinical and research applications. Recently, some studies have suggested to perform the segmentation based directly on the dMRI signal [194, 195, 196, 192]. One study demonstrated the feasibility of segmenting the corticospinal tract (CST) from the dMRI data for neonatal subjects in the dHCP dataset [192]. Another work segmented the 72 tracts from the TractSeg study [196]. It showed that it was possible to achieve a similar level of segmentation accuracy as TractSeg while using only six measurements from a single shell.
  • While most studies have focused on deep white matter tracts, a few studies have addressed the arguably more challenging task of segmenting the superficial white matter tracts, also known as the U-fibers [228, 187, 229, 230]. These tracts are critical for assessing brain connectivity as they account for the majority of the cortico-cortical connections. However, they are difficult to study due to their small size, partial volume effects, difficulty of reconstructing them with standard tractography techniques, as well as higher inter-subject variability [231, 230, 229]. Xue et al. [187, 232] developed a neural network model to segment 198 superficial white matter tracts via clustering the streamlines using a deep neural network. Streamlines were represented as point clouds and a contrastive learning approach was used to train the network. The authors reported remarkable accuracy (cluster identification rates of 87-99%, except for a neonatal dataset) for different age groups and health conditions including brain tumor patients. Another study [229] developed a hierarchical clustering method to identify representative groups of short association fibers from a population of subjects. This was accomplished by identifying fibers that were present in the majority of the subjects in a common (Talairach) space. The method was used to develop an atlas of 93 reproducible tracts, which could be then used to segment these tracts from whole-brain tractograms of test subjects. Another study suggested clustering of superficial white matter tracts based on their connectivity pattern using cortical brain parcellations inferred with an atlas [230].
@@
TABLE VII: A summary of selected non-tractography-based methods for automatic delineation/segmentation of white matter tracts.
 Reference
Model
Input
Number of tracts
Main findings
 [191]
atlas registration
diffusion tensor
11
The method performed comparably with manual tractography-based approach. It also worked well on an MS brain.
[192]
CNN
b0 image
1
This work only addresses the corticospinal tract (CST).
[163]
CNN
raw diffusion data
12
The method showed high test-retest reproducibility and generalizability across scanners.
[186]
CNN
orientation of the major FOD peaks
72
Modeling the label correlations improved the segmentation accuracy, especially for hard-to-segment tracts.
[193]
CNN
orientation of the major FOD peaks
72
Self-supervised pre-training using tractography data or multi-atlas segmentation data improved segmentation accuracy.
[194, 195]
CNN
dMRI data represented in a spherical harmonics basis
72
@@
This work evaluates the accuracy of the dMRI-based segmentation in terms of its overlap with positive and negative of responses of navigated transcranial magnetic stimulation.
[196]
CNN
dMRI data
72
@@
This work showed that it was possible to segment all 72 tracts form the TractSeg study using only six dMRI measurements as input.
[197]
CNN
diffusion tensor
25
@@
This method produced highly accurate and reproducible tract segmentations and it was generalizable to unseen datasets.
[198]
CNN
orientation of the major FOD peaks
72
@@
Only a single manually labeled image was used for training. A registration-base label propagation method was employed to synthesize additional labeled images.
II-E dMRI registration and tract-specific analysis
@@
Tract-specific studies, as their name implies, analyze dMRI-computed biomarkers on specific white matter tracts [233]. They can be of two types: (i) longitudinal studies consider the same subject(s) over time, while (ii) population studies compare different cohorts of subjects such as diseased versus control groups. They are among the most common and most fruitful types of dMRI studies because many neurodevelopmental deficits and neurological disorders are linked to damage to the tissue micro-structure on specific tracts [234, 235]. The success of these studies depends on ensuring that the same tracts are compared between scans/subjects, which is typically achieved via precise alignment of different brain scans. Regardless of the approach taken, this is a complex and error-prone computation. Existing computational methods vary greatly in terms of spatial accuracy and the required manual work [233]. Voxel-based morphometry methods are simple but incapable of accurate tract-specific comparisons [236], while semi-automatic and automatic tractography-based methods are limited by the tractography errors [237, 172]. Among the automatic methods, TBSS [78] is the most widely used. However, it suffers from a host of inaccuracies and shortcomings that have been pointed out by many studies in the past two decades [5, 238, 239]. Machine learning techniques have a unique potential to develop more reliable methods in order to address these limitations.
@@
A main source of error in many tract-specific analysis methods is due to poor registration. Some methods, such as TBSS, perform the registration based on FA images, which ignores the orientation-dependent micro-structure information [240, 239]. Leveraging the fiber orientation information can significantly aid in accurate alignment [241, 242]. Several deep learning-based dMRI registration methods have been proposed in the past few years and they have shown a tremendous potential for accurate and robust registration [243, 244, 245, 246]. Overall, these methods follow the recent advancements in deep learning-based registration, which are mainly based on fully convolutional networks, spatial transformer networks, and unsupervised training [247]. Additionally, to properly re-orient the white matter structures based on non-linear deformations they usually employ the finite strain technique [248]. One study has shown that using the diffusion tensor images in addition to anatomical (T1 and T2) images can lead to superior registration accuracy [243]. A subsequent work has shown that attention maps can be learned to weight the contribution of the two modalities (i.e., anatomical and micro-structural images) to achieve higher registration accuracy [246]. Another study used the tract orientation maps [37] for 40 white matter tracts as well as FA maps (for the moving and fixed image pairs) to compute 41 separate deformation fields [244]. A fusion neural network combined these into a final deformation field. On data from different age groups, scanners, and imaging protocols, this method was more accurate than existing DTI- and FOD-based registration techniques.
@@
A few studies have moved one step ahead and proposed novel computational methods to enable cross-subject tract-specific analysis. These methods perform both segmentation of individual tracts and alignment of data from different subjects/scans either directly or indirectly. Such methods are urgently needed since existing solutions such as TBSS are known to suffer from important limitations. We briefly discuss some of these methods below by separating them into those that use tractography and those that are based on deep learning.
@@
  • Tractography-based method. Prasad et al. [249] developed a framework based on streamline clustering, skeletonization based on highest streamline density, and registration based on geodesic curve matching. They showed that this method had a higher sensitivity than TBSS. A rather similar approach, based on multi-atlas clustering of streamlines with label fusion, was proposed and validated in [250]. Zhang et al. [251] used group-wise registration and analysis of whole-brain tractograms to compute an atlas of valid streamlines, which they then used to extract subject-specific fiber tracts. They used this method to analyze the brain micro-structure in terms of FA and MD for a population of autism spectrum disorder children and healthy controls. Deep learning methods are expected to perform these segmentation and registration operations with superior accuracy and speed.
@@
  • Deep learning-based methods. Segis-Net [252, 253, 254] is a novel deep learning method that performs a joint tract segmentation and registration. The segmentation module and the registration module are optimized jointly. The segmentation module uses diffusion tensor as input, while the registration module uses FA images as input. The loss function includes separate terms for segmentation and registration as well as a joint term that encourages tract segmentation overlap between pairs of registered images. The advantage of this approach is the joint learning of the segmentation and registration tasks and the reliance on an implicit unbiased reference image [254]. Segis-Net was applied to analyze six tracts in a longitudinal study including 8045 scans and was shown to lead to higher reproducibility than a non-learning method and a machine learning pipeline that tackled segmentation and registration tasks separately. Segis-Net also reduced the sample size requirements for tract-specific studies by a factor of up to three and it was orders of magnitude faster than methods based on conventional segmentation and registration techniques. A more recent method is TBSS++, where the authors develop an accurate and detailed dMRI atlas that includes 72 tracts [255]. For each test subject, the method segments individual tracts and registers the subject data onto the atlas space. This method offers higher reproducibility and better robustness to measurement noise than TBSS. In a different work, an autoencoder has been proposed to detect abnormal micro-structure indices on white matter tracts [256]. The autoencoder input consists of micro-structural indices on white matter tract profiles. Following the standard autoencoder approach, the model is trained on a population of normal subjects to learn to reconstruct healthy brain data. When applied on data from pathological brains, the autoencoder can detect abnormal micro-structure values via large reconstruction errors. The authors validate this method on data from patients with various neurological disorders including epilepsy and schizophrenia. These experiments show that the new method is more accurate than standard abnormality detection techniques. The authors characterize their method as moving beyond group-wise analysis and enabling assessment of single patients at an individual level. However, one can argue that techniques such as TBSS++ can offer the same capability.
II-F Segmentation

### v2

@@ IV-G2 Methods that avoid tractography
In addition to the streamline similarity measures and anatomical information described above, tractography-based methods have resorted to various other information to improve their accuracy. These include homology between the brain hemispheres, spatial and shape priors, and population averages (atlases). However, all tractography-based methods are inherently limited by the errors in streamline tracing [37, 302]. Moreover, they typically involve several processing steps such as streamline propagation, white matter segmentation, gray matter parcellation, and clustering or similar computations to extract individual tracts. Each of these computations may introduce its own errors that depend on method settings, and it is difficult to jointly optimize all these processing steps. Moreover, some of these computations, e.g., tractography and streamline clustering, can require long computation times.
@@
TABLE VIII: A summary of selected non-tractography-based methods for automatic delineation/segmentation of white matter tracts.
  Reference Model Input Number of tracts Methodology and results
  TractSeg [17] Set of 2D CNNs Orientation of the major FOD peaks 72 TractSeg was the first work to demonstrably show that deep learning models can segment various white matter tracts without the need for tractography. It achieved an average Dice Similarity Coefficient of 0.84 on HCP data and 0.82 on clinical quality data.
Liu et al. [303] CNN Orientation of the major FOD peaks 72 This work showed that modeling the label correlations improved the segmentation accuracy, especially for tracts that were more difficult to segment. Experiments show that this method works better than an atlas-based method and TractSeg [17].
Lu et al. [304] CNN Orientation of the major FOD peaks 72 This work advocates for self-supervised pre-training of the CNN using two pretext tasks: (1) tractography density prediction, (2) segmentation with noisy labels generated with an atlas-based method. It reports improved segmentation accuracy.
@@
Lucena et al. [305, 306] CNN dMRI data represented in a spherical harmonics basis 72 This work evaluates the accuracy of the dMRI-based segmentation in terms of its overlap with the responses of navigated transcranial magnetic stimulation.
@@
Kebiri et al. [307] CNN dMRI data 72 This work showed that it was possible to segment all 72 tracts form the TractSeg study using only six dMRI measurements as input. However, only one independent dataset was used for validation.
HAMLET [243] CNNs with rotation equivariance Raw diffusion data 12 The CNN model uses rotationally equivariant convolutional operations, thereby eliminating the need to learn data orientations. Experiments show high test-retest reproducibility and generalizability on low-quality data from an external scanner.
@@
Neuro4Neuro [308] CNN Diffusion tensor 25 This method produced highly accurate and reproducible tract segmentations and it was generalizable to unseen data from external scanners and from dementia patients.
@@
Xu et al. [309] CNN Orientation of the major FOD peaks 72 This work used a registration-base label propagation method to synthesize noisy labels for unlabeled images, thereby increasing the size of labeled training data. The authors reported that with this approach, only a single manually labeled image was sufficient for training.
To avoid these errors and limitations, several studies have proposed to segment the tracts on diffusion tensor or fiber orientation images, thereby avoiding the tractography. Some of the classical machine learning methods that have been used for this purpose include Markov Random Fields [310], k-nearest neighbors technique [311], level-set methods [312, 313, 314], template/atlas-based techniques [315, 302].
More recently, deep learning has shown unprecedented accuracy for this application as well [17, 316]. Wasserthal et al. [54] represented overlapping/crossing tracts using a simple but powerful concept that they named tract orientation maps (TOMs). Each tract is represented by a separate TOM, where each voxel contains a vector representing the local orientation of that tract. If a tract does not cross a certain voxel, the value of that voxel in the TOM for that tract will be a vector of zeros. In their framework, a standard method such as CSD is applied to compute the FOD and the FOD peaks are extracted to build the TOMs. These peak orientations are fed into an FCN, which is trained to compute the TOMs. The TOMs can be used as a prior or direct input for bundle-specific tractography. The authors claim that their method is a viable solution to excessive false positive rate of standard tractography techniques. They support this claim by demonstrating that their method leads to superior tract bundle reconstruction compared with several state of the art methods such as TractQuerier [291], RecoBundles [279], and WhiteMatterAnalysis (WMA) [271]. In subsequent works, they build upon this method and extend it in a few important directions. In TractSeg [17], they develop an FCN-based method to map the TOMs to the segmentation probability maps for each tract. To enable the method to run on limited GPU memory while working on full-resolution images to segment the complete set of 72 tracts, their model is a 2D FCN that works on axial, coronal, and sagittal slices separately. TractSeg achieves a mean Dice Similarity Coefficient (DSC) of 0.84. On all tracts except for commissure anterior and fornix it achieves a mean DSC of higher than 0.75. While the tract segmentation masks generated by TractSeg can be used for accurate bundle-specific tractography [17], a more complete method is presented in a subsequent work by the same authors [317], where a separate deep learning model segments the start and end regions of each tract.
A number of studies have built upon the methods proposed by Wasserthal et al. [54, 317]. A representation similar to TOM was proposed in HAMLET [243], where the model output is a tensor field that indicates the presence and local orientation of a tract. Li et al. [308] trained a CNN to segment 25 white matter tracts directly from diffusion tensor images. Their extensive experiments showed high segmentation accuracy, low scan-rescan variability in tract-specific FA analysis, good generalizability to cross-center data, and successful application in a tract-specific population study on dementia patients. Their method was more accurate than tractography-based techniques, while being orders of magnitude faster. They also found that adding anatomical (T1-weighted) images to the diffusion tensor as input did not improve the segmentation accuracy. Moreover, using the directions of the three major peaks led to lower accuracy than diffusion tensor [308]. Liu et al. followed an approach similar to TractSeg but proposed to model the correlation between tract labels [303]. Specifically, they introduced an auxiliary neural network that mapped the native tract labels into a lower-dimensional label space. The main segmentation network predicted this lower-dimensional label, which could be used to predict the tract segmentation maps. The authors argue that this approach makes the task of tract segmentation simpler to learn and show that it can improve the segmentation of certain tracts such as the fornix that are especially difficult to segment with TractSeg. Similarly, Lu et al. follow the general approach proposed by TractSeg, but propose self-supervised pre-training [304] and transfer learning [318] to reduce the required labeled data. These works have shown that as few as five labeled images may be sufficient to train CNNs for segmenting certain tracts. To reduce the required manual labelling, another study proposed a deep learning registration method to align the labeled images to unlabeled images, thereby synthesizing a large training dataset [309]. Experiments showed that even a single labeled image was sufficient to train an accurate deep learning model.
In order to determine the orientation of major fascicles, previous works have relied on computation of the diffusion tensor or FOD. However, none of these intermediate computations have an unambiguous biophysical meaning and they entail unavoidable estimation errors. Moreover, the intermediate computations for most existing methods assume availability of dense multi-shell measurements, which are not acquired in many clinical and research applications. Recently, some studies have suggested to perform the segmentation based directly on the dMRI signal [305, 306, 307, 319]. One study demonstrated the feasibility of segmenting the corticospinal tract (CST) from the dMRI data for neonatal subjects in the dHCP dataset [319]. Another work segmented the 72 tracts from the TractSeg study [307]. It showed that it was possible to achieve a similar level of segmentation accuracy as TractSeg while using only six measurements.
IV-G3 Superficial white matter analysis methods
While most studies have focused on deep white matter, a few studies have addressed the arguably more challenging task of segmenting the superficial white matter tracts, also known as the U-fibers [320, 275, 321, 322]. These tracts are critical for assessing brain connectivity as they account for the majority of the cortico-cortical connections. However, they are difficult to study with standard tractography techniques due to their small size, partial volume effects, and high inter-subject variability [323, 322, 321]. Xue et al. [275] developed a neural network model to segment 198 superficial white matter tracts via clustering the streamlines using a deep neural network. Streamlines were represented as point clouds and a contrastive learning approach was used to train the network. The authors reported remarkable accuracy (cluster identification rates of 87-99%, except for a neonatal dataset) for different age groups and health conditions including brain tumor patients. Another study [321] developed a hierarchical clustering method to identify representative groups of short association fibers from a population of subjects. This was accomplished by identifying fibers that were present in the majority of the subjects in a common (Talairach) space. The method was used to develop an atlas of 93 reproducible tracts, which could then be used to segment these tracts from whole-brain tractograms of test subjects. Another study suggested clustering of superficial white matter tracts based on their connectivity patterns using cortical brain parcellations inferred with an atlas [322].
IV-H dMRI registration
@@ IV-G3 Superficial white matter analysis methods
Tract-specific studies, as their name implies, study the organization, development, and tissue microstructure of specific white matter tracts [324]. They can be of two types: (i) longitudinal studies consider the same subject(s) over time, while (ii) population studies compare different cohorts of subjects such as diseased versus control groups. They are among the most common dMRI studies because many neurodevelopmental deficits and neurological disorders are linked to damage to the tissue microstructure on specific tracts [325, 326]. The success of these studies depends on ensuring that the same tracts are compared between scans/subjects, which is typically achieved via precise alignment of different brain scans. Regardless of the approach taken, this is a complex and error-prone computation. Existing computational methods vary greatly in terms of spatial accuracy and the required manual input [324]. Voxel-based morphometry methods are simple but incapable of accurate tract-specific comparisons [327], while semi-automatic and automatic tractography-based methods are limited by the tractography errors [328, 264]. Among the automatic methods, TBSS [174] is the most widely used. However, it suffers from important inaccuracies and shortcomings [14, 329, 330]. Machine learning techniques have a unique potential to develop more reliable methods to address these limitations.
@@
A main source of error in many tract-specific analysis methods is poor registration. Some methods, such as TBSS, perform the registration based on FA images, which ignores the orientation-dependent microstructure information. Leveraging the fiber orientation information may significantly improve alignment accuracy [331, 332]. Several deep learning-based dMRI registration methods have been proposed in the past few years and they have shown a tremendous potential for accurate and robust registration [333, 334, 335, 336]. Overall, these methods follow the recent advancements in deep learning-based registration, which are mainly based on fully convolutional networks, spatial transformer networks, and unsupervised training [337]. Additionally, they usually use the finite strain technique [338] to properly reorient the white matter structures based on the computed non-linear deformations. One study has shown that using the diffusion tensor images in addition to anatomical (T1- and T2-weighted) images can lead to higher registration accuracy [333]. A subsequent work has shown that attention maps can be learned to weight the contribution of the two modalities (i.e., anatomical and microstructural images) to achieve better registration [336]. Another study used the tract orientation maps [54] for 40 white matter tracts as well as FA maps to compute 41 separate deformation fields [334]. A fusion neural network combined these into a final deformation field. On data from different age groups, scanners, and imaging protocols, this method was more accurate than existing DTI- and FOD-based registration techniques.
IV-I Tract-specific analysis
@@
A few studies have proposed machine learning methods for cross-subject tract-specific analysis. Most these methods, but not all, are based on tract segmentation and alignment of data from different subjects/scans. Such methods are urgently needed since existing solutions such as TBSS are known to suffer from important limitations. These works are briefly discussed below under two classes of approaches.
IV-I1 Tractography-based methods
@@
Prasad et al. [339] developed a framework that consisted of streamline clustering, skeletonization based on highest streamline density, and registration based on geodesic curve matching. They showed that this method had a higher sensitivity than TBSS. Jin et al. proposed a similar approach based on multi-atlas clustering of streamlines with label fusion [340]. Zhang et al. [341] used group-wise registration and analysis of whole-brain tractograms to compute an atlas of valid streamlines, which they then used to extract subject-specific fiber tracts. They used this method to analyze the tissue microstructure in terms of FA and MD for a population of autism spectrum disorder children and healthy controls. If the goal is to predict subject-level neurocognitive scores, cross-subject data alignment or spatial correspondence can be side-stepped. TractGeoNet [342], for example, uses a neural network to directly predict the score from streamline-level tissue microstructure information.
IV-I2 Deep learning-based methods
@@
Segis-Net [23] is a deep learning method for tract segmentation and registration. The segmentation and registration modules are optimized jointly. The segmentation module uses diffusion tensor as input, while the registration module uses FA images as input. The loss function includes separate terms for segmentation and registration as well as a joint term to encourage tract segmentation overlap between pairs of registered images. The advantage of this approach is the joint learning of the segmentation and registration tasks and reliance on an implicit unbiased reference image [343]. Segis-Net was applied to analyze six tracts in a longitudinal study including 8045 scans and was shown to lead to higher reproducibility than a non-learning method and a machine learning pipeline that addressed segmentation and registration tasks separately. Segis-Net also reduced the sample size requirements for tract-specific studies by a factor of up to three and it was orders of magnitude faster than methods based on conventional segmentation and registration techniques. In another work, the authors develop an accurate and detailed dMRI atlas that includes 72 tracts [344]. For each test subject, the method segments the tracts and registers the subject data onto the atlas. This method offers higher reproducibility and better robustness to measurement noise than TBSS. In a different work, an autoencoder has been proposed to detect abnormal tissue microstructure indices on white matter tracts [345]. The autoencoder input consists of microstructural indices on white matter tract profiles. Following the standard autoencoder approach, the model is trained on a population of normal subjects to learn to reconstruct healthy brain data. When applied on data from pathological brains, the autoencoder can detect abnormal microstructure indices via large reconstruction errors. The authors validate this method on data from patients with various neurological disorders including epilepsy and schizophrenia. These experiments show that the new method is more accurate than standard abnormality detection techniques. The authors characterize their method as moving beyond group-wise analysis and enabling assessment of single patients at an individual level. However, one can argue that techniques such as [23, 344] can offer the same capability.
IV-J Segmentation